OpenStudy (anonymous):

Nestor borrows $10,000 from the shady loan shark, Mr. Wildcat. Mr. Wildcat tells him he must repay the loan in 1 month and pay an additional $1000 in interest. If Mr. Wildcat is charging simple interest on his loan, what is the annual interest rate?

OpenStudy (anonymous):

Do I use the formula A=P(1+rt)

OpenStudy (tkhunny):

Part of it. You have this: P = 10000 t = 1/12 A = 10000 + 1000 = 11000 You have only to solve for 'r'.

OpenStudy (anonymous):

ok awesome! that is what I thought!

OpenStudy (anonymous):

so would my answer be 120%

OpenStudy (tkhunny):

That's why it's "Shady Loan Shark" and not "Polite and Legal Banker"! Yikes.

OpenStudy (anonymous):

lol thanks!